what is the formula of emi?

Grade:12th Pass

1 Answers

Sai Kiran
11 years ago

HI..

This is the formula to caluculate:
E = P×r×(1 + r)n/((1 + r)n - 1)
E is EMI
where P is Priniple Loan Amount
r is rate of interest calualted in monthly basis it should be = Rate of Annual interest/12/100
if its 10% annual ,then its 10/12/100=0.00833
n is tenture in number of months
Eg: For 100000 at 10% annual interest for a period of 12 months
it comes  to 100000*0.00833*(1 + 0.00833)12/((1 + 0.00833)12 - 1) = 8792

HOPE THIS HELPS.. :D